OpenStudy (anonymous):

HELPPPPPPPPP Use the simple interest formula to find the missing value. Rates are annual rates. Assume 360 days per year and 30 days per month. p = ?, r = 4.5%, t = 10 days, i = $26.68 A. p = $21,344.00 B. p = $711.47 C. p = $7114.67 D. p = $21,670.09

OpenStudy (anonymous):

Simple interest Interset formula is given as follows: I =prt => P = I/RT where I = interest accrued p = principal r = rate per year Given I = 26.68 t= 10 days I= 4.5% per year. 4.5% is the rate per year: => Rate per day is 4.5%/(360) = 0.045/360 Therefore P = 26.68/[(0.045/360)] *10 = (26.68 8 360)/(0.045*10) =21344 Therefore the answer is A A. p = $21,344.00